Can someone show me how to write the formula for this in math terms and solve it? I tried several ways but could not write the formula or solve it correctly. Thanks in advance for your time and efforts.
Pat is 3 times as old as she was 14 years ago. In 7 years, she will be 4 times as old as she was 14 years ago. How old is she now?

Answer by ankor@dixie-net.com(22740)

Pat is 3 times as old as she was 14 years ago. In 7 years, she will be 4 times as old as she was 14 years ago.
How old is she now?
:
Let a = Pat's age now
:
Write an equation for each statement, simplify:
:
"Pat is 3 times as old as she was 14 years ago."
a = 3(a-14)
a = 3a - 42
42 = 3a - a
42 = 2a
a = 42/2
a = 21 yrs now
:
" In 7 years, she will be 4 times as old as she was 14 years ago.'
a + 7 = 4(a-14)
a + 7 = 4a - 56
a = 4a - 56 - 7
a = 4a - 63
63 = 4a - a
63 = 3a
a = 63/3
a = 21 yrs now, which we already knew.